The perimeter of a rectangle is 20 inches. The area of the same rectangle in 18 square inches. Find the width of the rectangle to the nearest tenth of an inch.
Answer by rfer(16322)

2x+2y=20
xy=18
--------------
2x+20-2y
x=10-y
-------------
(10-y)(y)=18
-y^2+10y-18=0
y^2-10y+18=0
----------------
y=(-b+or- sq rt b^2-4ac)/2a
y=(-10+or- 5.29)/2
y=15.29/2
y=7.645
---------------
7.645x=18
x=2.355
-------------
2.355*7.645=18
--------------
2(2.355)+2(7.645)=20